OpenStudy (anonymous):

What value of w does the system have no solution: {3A + 5B = -12 wA - 15B = 6

OpenStudy (solomonzelman):

you want to get it to the point where the left side of the second equation, is a multiple of the left side of the first equation.

OpenStudy (solomonzelman):

Like if your problem had a system of: \(\large\color{black}{ 12Z+10Y=7 \\kZ -5Y=5 }\) then: \(\large\color{black}{ 12Z+10Y=7 \\-2(k)Z +10Y=-10 }\) \(\large\color{black}{ -2(k)=12 }\) \(\large\color{black}{ k=-6 }\)

OpenStudy (solomonzelman):

And there, we will have when k=-6, \(\large\color{black}{ 12Z+10Y=7 \\12Z +10Y=-10 }\)

OpenStudy (solomonzelman):

and you see that THIS has no solution, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Could you help me with this problem? Because i am confused...

OpenStudy (solomonzelman):

I am trying, but I keep getting disconnected.

OpenStudy (anonymous):

gotcha

OpenStudy (solomonzelman):

\(\large\color{black}{ 3A + 5B = -12 \\ wA - 15B = 6 }\) I'll set you up, lets divide the second equation by -3, we get: \(\large\color{black}{ 3A + 5B = -12 \\ (-\frac{1}{3})wA + 5B = -2 }\)

OpenStudy (solomonzelman):

you now need to a value of w, for which (1/3)(w) is equal to 3

OpenStudy (solomonzelman):

\(\large\color{black}{ (-\frac{1}{3})w=3 }\), then w=?

OpenStudy (anonymous):

um...

OpenStudy (anonymous):

I'm not sure

OpenStudy (solomonzelman):

\(\large\color{black}{ (-\frac{1}{3})w=3 }\) multiply both sides times -3, and tell me what w equals.

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

w = -9

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

So now what?

OpenStudy (solomonzelman):

so, in order to make your system of equations, \(\large\color{black}{ 3A + 5B = -12\\ wA - 15B = 63 }\) have no solutions, \(\large\color{black}{w=-9 }\)

OpenStudy (solomonzelman):

you are done:) this is it,.

OpenStudy (anonymous):

okay! Cool! Thanks.

OpenStudy (solomonzelman):

Not a problem!

OpenStudy (anonymous):

Can you help me with this problem? Please?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Let's see, I'll apload it, and do some labelling part:

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/yx3vavvpb3